OpenStudy (anonymous):

Q.1 The survivor of a shipwreck lands on an island which is 300m from a vertical cliff. He sees a ship anchored between the island and the cliff. A blast from the ship's horn is heard twice with a time lapse of 4s. Calculate the distance d of the ship from the island (Assume that the speed of the sound=330ms). Q.2 A man stands between two cliffs and claps his hands once he hears two echoes, one after 0.5s and the other 0.3s after the first. If the speed of sound is 330ms what is the distance between the two cliffs?

OpenStudy (anonymous):

@emcrazy14 , where are you facing problem??

OpenStudy (anonymous):

Actually, I can't solve the first question. I solved the second one but I'm not sure if the answer is correct? Will it be 214.5m? And can you pleaase show me how to do the first one?

OpenStudy (mos1635):

 t1=x/V t2=(300-x)/v Δt=(2t2+t1)-t1 Δt=4